26.8 g Na2SO4.NH2O has 12.6 g water .Find the value of n

Grade:

1 Answers

Vikas TU
4 years ago
mass= 26.8-12.6=14.2
molar mass= (23x2)+(32+64)=142
 
there its moles =14.2/142
                          =0.1
there ration of moles of Na2SO4:H2O =0.1:0.7
hence the whole number ration =    1:7 (meaning there are 7 molecules of water in this hydrated salt)
hence n=7